you pay $105 for eight tickets. If student tickets cost $10 and adult tickets cost $5, how many of each type did you buy?
Answer by checkley79(3341)

10x+5y=105
x+y=8 or x=8-y
10(8-y)+5y=105
80-10y+5y=105
-5y=105-80
-5y=25
y=-25/5
y=-5
This tells me one of the factors you have are incorrect.
Test:
All tickets are adults.
8*10=$80 not enough money.